OpenStudy (anonymous):

int x=10; x+=5; gives the value of x as 5,15,6 or 16?

OpenStudy (anonymous):

13 Because, x=10, given. So, x+5= 10+5= 15 (Just plug in the value 10 in place of x)

OpenStudy (anonymous):

Thanks a lot but I missed something in that question and the question is: int x =10; x+=5; gives the value of x as there is four options for answer 5,15,6 or 16.